OpenStudy (osuxbucks15):

The equation (x + 6)2 + (y + 4)2 = 36 models the position and range of the source of a radio signal. Describe the position of the source and the range of the signals.

OpenStudy (osuxbucks15):

@radar can u help me please?

OpenStudy (radar):

You are looking at a circular coverage with a range of 6 units and a location at point P(-6,-4) See the formula for a circle whose center is located away from the origin.

OpenStudy (osuxbucks15):

(x + 6)2 + (y + 4)2 = 36 this one?

OpenStudy (radar):

Yes, that is a circle whose radius is sq rt of 36. and center at (-6,-4)

OpenStudy (radar):

See "Formula for circle"

OpenStudy (radar):

That is the formula for area of circle. The formula for a circle whose center is located at the origin is:\[x ^{2} + y ^{2}= r ^{2}\]

OpenStudy (osuxbucks15):

o ok

OpenStudy (radar):

For circles whose center is offset from the origin, the formula is slightly different, that is why I want you to look it up.

OpenStudy (osuxbucks15):

o ok so i should look up "Formula for circle" on google

OpenStudy (osuxbucks15):

(x-h)2 + (y-k)2 = r2 what about this

OpenStudy (radar):

To save your valuable time, the formula for a circles whose center is offset from the origin is:\[(x-h)^{2}+(y-k)^{2}=r ^{2}\]

OpenStudy (osuxbucks15):

ok so which numbers do i plug in for x,y and k

OpenStudy (radar):

That is what I was typing as you posted, which proves that you can type faster than I lol

OpenStudy (osuxbucks15):

no I'm an alright typer i copied and pasted that!

OpenStudy (osuxbucks15):

ok now back to the question

OpenStudy (radar):

Look at your formula (x+6)^2 + (y+4)^2=36 the answer to your last question resides in that formula. what is the -h? what is the -k? what is r^2?

OpenStudy (osuxbucks15):

ok h is 6 4 is k and r=36

OpenStudy (radar):

-h = 6 What does h =?

OpenStudy (radar):

You are rushing. r^2 = 36, what is r?

OpenStudy (osuxbucks15):

h=-6? and 18=r

OpenStudy (radar):

\[r ^{2}=36\] What is r?

OpenStudy (osuxbucks15):

wouldn't it be 36/2 to get r

OpenStudy (radar):

and don't forget to solve for k when -k=4

OpenStudy (osuxbucks15):

k=-4

OpenStudy (radar):

r SQUARED =36 /does 18 squared = 36??? think about it.

OpenStudy (osuxbucks15):

r=6

OpenStudy (radar):

Yes, do you understand the radio coverage?

OpenStudy (osuxbucks15):

is that like the squared thing that we just did

OpenStudy (osuxbucks15):

because if it is not that then i don't no it

OpenStudy (radar):

 It is a circular coverage with a radius of 6 (as we don't know if it is miles or km or what, the problem did not say so we can only say 6 units. Whose center is located at P(-6,-4) again we don't know the distance units.